Use the elimination method to solve the given system of equations. Please show all work
8x + 11y = 20
5x - 11y = -59

Answer by mananth(16946)

8 x + 11 y = 20 .............1
Total value
5 x -11 y = -59 .............2
Eliminate y
Add the two equations
13 x = -39
/ 13
x = -3
plug value of x in (1)
8 x + 11 y = 20
-24 + 11 y = 20
11 y = 20 + 24
11 y = 44
y = 4
x=-3, y=4